# Efficient and Workload-Aware LLM Serving via Runtime Layer Swapping and KV Cache Resizing

### v1

@@ B.2  Greedy Selection Policy
Profiling Setting. Following the setup in [37, 57, 44], we use a small calibration subset from the WikiText2 dataset, with the sequence length of 2,048. While LIS incorporates static layer-level metrics and model-level output feedback, its design avoids overfitting the specific calibration set. Once computed, the LIS ranking is fixed and reused across deployments, requiring no online re-tuning.
@@
B.3  Evaluation
To evaluate the effectiveness and generalizability of the LIS-based layer selection strategy, we compare it against several ordering baselines using perplexity across four models: Vicuna [42], Llama 2 [64], Llama 3 [22], and CodeLlama [56]. The comparison includes the following baselines: Front-to-Back—layers are swapped sequentially from the input (first layer) to the output (last layer); Back-to-Front—the reverse order, starting from the final layer and proceeding backward; Random—a randomly shuffled layer order, averaged over multiple runs to reduce variance.
@@ C  Implementation Details
Implementation. MorphServe is built on top of SwiftLLM [58, 31], with approximately 2,200 lines of Python and 500 lines of C++/CUDA. It adds runtime support for dynamic layer swapping and elastic KVC resizing with minimal changes to the scheduler and attention mechanisms, such as FlashAttention [12, 11] and PagedAttention [33], and remains compatible with state-of-the-art LLM inference engines such as vLLM [33]. At initialization, full-precision and quantized transformer layer weights (FP16, W8, W4) are preloaded into pinned CPU memory, and all GEMM kernels are precompiled using dummy data to eliminate runtime compilation overhead. GPU memory regions for each layer are preallocated, enabling in-place weight swapping via cudaMemcpyAsync without pointer remapping. For KV cache resizing, we extend PagedAttention to support block-level reallocation and remapping through dynamic memory registration. Morphing and decoding are executed on separate CUDA streams to ensure efficient asynchronization and minimize interference with token generation.
@@
Serving Traces. The Azure LLM Inference Dataset 2023 trace [4, 51] is a publicly released dataset capturing anonymized LLM request logs from Azure’s cloud infrastructure. It includes request arrival times, prompt, and output lengths statistics. The dataset is publicly available at https://github.com/Azure/AzurePublicDataset. For our evaluation, we sample 72 seconds of traffic with a downscaling factor of 4.75×4.75$\times$4.75 × to match the hardware memory footprint and enable simulation of large-batch request bursts. BurstGPT [67] is a real-world LLM inference workload trace collected from a university campus. It captures naturally occurring burst patterns resulting from student and faculty interactions with deployed chatbots and LLM-based tools. The trace includes detailed request metadata such as arrival timestamps, prompt lengths, and session-level characteristics, enabling realistic simulation of latency-sensitive serving conditions. It is publicly available at https://github.com/HPMLL/BurstGPT. For our evaluation, we also extract a 72-second segment and apply a 1.75×1.75$\times$1.75 × downscaling factor to simulate saturation-level conditions. This trace is used to benchmark MorphServe’s responsiveness and adaptation under real-world burst traffic.
Evaluation Datasets. GovReport [27] is a long-form summarization dataset consisting of U.S. government reports paired with expert-written summaries. It is publicly available at https://huggingface.co/datasets/launch/gov_report. We use it to benchmark summarization quality and stress-test long input handling. Average document length exceeds 2,000 tokens, making it suitable for evaluating memory-intensive generation. QMSum [77] is a query-based meeting summarization dataset comprising multi-party meeting transcripts with user-specified queries and corresponding abstractive summaries. Available at https://github.com/Yale-LILY/QMSum, it tests both summarization and task-oriented comprehension under long-context inputs. DuReader [23] is a Chinese machine reading comprehension dataset with over human-annotated question-answer pairs from Baidu search logs. It covers open-domain QA with a range of answer formats. We use the English-translated version and evaluate factual correctness. The dataset is hosted at https://github.com/baidu/DuReader . Multi-News [15] is a multi-document summarization dataset containing news articles from multiple sources clustered around the same event, with human-written summaries. It is accessible at https://github.com/Alex-Fabbri/Multi-News. This dataset evaluates the model’s ability to synthesize content across multiple documents and is especially useful for benchmarking performance on broad-context summarization.
@@ D  Limitations and Broader Impacts
Limitations. While MorphServe is practical and effective for dynamic LLM serving, several limitations remain. To support runtime layer swapping, MorphServe stores both full-precision and quantized variants in host memory. Although this increases memory usage, the overhead is typically under 2×2$\times$2 × the model size and is well accommodated by modern LLM serving clusters. Future work may further reduce this cost by streaming layers from SSD to host memory on demand or directly fetching them from SSD via GPUDirect Storage (GDS) [48]. MorphServe currently applies morphing at the transformer layer level. While effective, finer-grained adaptation, such as independently adjusting attention and MLP submodules, could unlock additional efficiency and precision flexibility. MorphServe reacts to system pressure in real time but does not anticipate upcoming surges. Integrating lightweight workload forecasting could enable proactive morphing decisions and further improve responsiveness under bursty traffic.

### v2

@@ A.2  Greedy Selection Policy
Profiling Setting. Following the setup in Lin et al. (2024a); Shao et al. (2023); Ma et al. (2024), we use a small calibration subset from the WikiText2 dataset, with the sequence length of 2,048. While LIS incorporates static layer-level metrics and model-level output feedback, its design avoids overfitting the specific calibration set. Once computed, the LIS ranking is fixed and reused across deployments, requiring no online re-tuning.
@@
A.3  Evaluation
To evaluate the effectiveness and generalizability of the LIS-based layer selection strategy, we compare it against several ordering baselines using perplexity across four models: Vicuna lmsys , Llama 2 Touvron et al. (2023), Llama 3 Grattafiori et al. (2024), and CodeLlama Roziere et al. (2023). The comparison includes the following baselines: Front-to-Back—layers are swapped sequentially from the input (first layer) to the output (last layer); Back-to-Front—the reverse order, starting from the final layer and proceeding backward; Random—a randomly shuffled layer order, averaged over multiple runs to reduce variance.
@@ A.3  Evaluation
B  Additional Experimental Results
B.1  Independence of Layer Quantization Effects
To demonstrate the independence of layer quantization effects, we conducted a direct validation using Llama 2 7B on the WikiText dataset. We progressively quantized the first $N$ layers ($N=0,1,2,4,8,16$) and measured the additional perplexity increase by further quantizing either layer 19 or layer 24. As shown in Table 7, the added perplexity from quantizing layer 19 remains consistently around 0.0035, regardless of how many earlier layers were already quantized. A similar pattern holds for layer 24 (0.0018-0.0020). This result provides support for the independence of layer-wise quantization effects.
Table 7: Evaluating independence of layer-wise quantization effects on perplexity under different quantization conditions.
Quantized Layers None First 1 Layer First 2 Layers First 4 First 8 First 16
PPL 5.472089 5.473293 5.474974 5.484205 5.503622 5.539723
PPL (+ layer 19 quant.) 5.475612 5.476802 5.478512 5.487739 5.507146 5.543575
Effect (layer 19) 0.003523 0.003509 0.003538 0.003533 0.003524 0.003852
PPL (+ layer 24 quant.) 5.473969 5.475141 5.476794 5.486121 5.505570 5.541744
Effect (layer 24) 0.001880 0.001848 0.001820 0.001916 0.001948 0.002021
B.2  A Verbatim Example for Mixed-Precision Serving
We further conducted a verbatim experiment to show MorphServe’s mixed-precision performance using Llama 3 8B Instruct, generating 128 output tokens under different serving scenarios with the following input prompt: “Zebras are primarily grazers and can subsist on lower-quality vegetation. They are preyed on mainly by lions, and typically flee when threatened but also bite and kick.”
  • Response (FP16): That’s correct! Zebras are indeed primarily grazers, and they are able to survive on a diet of lower-quality vegetation, such as grasses and shrubs. This is because they have a specialized digestive system…
  • Response (Mixed-serving case 1): That’s correct! Zebras are indeed primarily grazers (FP16), and they are well adapted to survive on a diet of grasses, leaves (W4), and other low-quality vegetation. Their digestive system is designed (FP16)…
  • Response (Mixed-serving case 2): That’s correct! Zebras are indeed primarily grazers (W4), and they are able to survive on a diet of lower-quality vegetation (FP16). Their digestive system is specially designed to break down and extract nutrients (W4)…
  • Response (W4): That’s correct! Zebras are indeed primarily grazers, and they are well adapted to survive on a diet of grasses, leaves, and other low-quality vegetation. Their digestive system is designed to break down…
MorphServe’s mixed-precision serving allocates compute budget adaptively, ensuring robust performance without compromising user experience.
B.3  Effectiveness Across Quantization Algorithms
Table 8 shows the performance of MorphServe with AWQ and Uniform INT4 quantization vs. FP16 on Vicuna 7B v1.5, using the QMSum dataset and Azure trace.
Table 8: Performance of MorphServe with AWQ and Uniform INT4 quantization vs. FP16 on Vicuna 7B v1.5, using the QMSum dataset and Azure trace.
Quantization (INT4) Metric Static Quantization MorphServe Full Precision
TTFT P95 (s) 0.4086 0.7281 6.2376
AWQ TPOT P99 (s) 0.0846 0.1123 0.1384
F1 Score 11.77 12.85 13.01
TTFT P95 (s) 0.3958 0.7034 6.2376
Uniform TPOT P99 (s) 0.0820 0.1095 0.1384
F1 Score 11.32 12.57 13.01
@@ C  Implementation Details
Implementation. MorphServe is built on top of SwiftLLM shengyu Liu and Jover Qian ; Jiang et al. (2024), with approximately 2,200 lines of Python and 500 lines of C++/CUDA. It adds runtime support for dynamic layer swapping and elastic KVC resizing with minimal changes to the scheduler and attention mechanisms, such as FlashAttention Dao et al. (2022); Dao (2023) and PagedAttention Kwon et al. (2023), and remains compatible with state-of-the-art LLM inference engines such as vLLM Kwon et al. (2023). At initialization, full-precision and quantized transformer layer weights (FP16, W8, W4) are preloaded into pinned CPU memory, and all GEMM kernels are precompiled using dummy data to eliminate runtime compilation overhead. GPU memory regions for each layer are preallocated, enabling in-place weight swapping via cudaMemcpyAsync without pointer remapping. For KV cache resizing, we extend PagedAttention to support block-level reallocation and remapping through dynamic memory registration. Morphing and decoding are executed on separate CUDA streams to ensure efficient asynchronization and minimize interference with token generation.
@@
Serving Traces. The Azure LLM Inference Dataset 2023 trace Azure (2024); Patel et al. (2024) is a publicly released dataset capturing anonymized LLM request logs from Azure’s cloud infrastructure. It includes request arrival times, prompt, and output lengths statistics. The dataset is publicly available at https://github.com/Azure/AzurePublicDataset. For our evaluation, we sample 72 seconds of traffic with a downscaling factor of $4.75\times$ to match the hardware memory footprint and enable simulation of large-batch request bursts. BurstGPT Wang et al. (2024) is a real-world LLM inference workload trace collected from a university campus. It captures naturally occurring burst patterns resulting from student and faculty interactions with deployed chatbots and LLM-based tools. The trace includes detailed request metadata such as arrival timestamps, prompt lengths, and session-level characteristics, enabling realistic simulation of latency-sensitive serving conditions. It is publicly available at https://github.com/HPMLL/BurstGPT. For our evaluation, we also extract a 72-second segment and apply a $1.75\times$ downscaling factor to simulate saturation-level conditions. This trace is used to benchmark MorphServe’s responsiveness and adaptation under real-world burst traffic.
Evaluation Datasets. GovReport Huang et al. (2021a) is a long-form summarization dataset consisting of U.S. government reports paired with expert-written summaries. It is publicly available at https://huggingface.co/datasets/launch/gov_report. We use it to benchmark summarization quality and stress-test long input handling. Average document length exceeds 2,000 tokens, making it suitable for evaluating memory-intensive generation. QMSum Zhong et al. (2021) is a query-based meeting summarization dataset comprising multi-party meeting transcripts with user-specified queries and corresponding abstractive summaries. Available at https://github.com/Yale-LILY/QMSum, it tests both summarization and task-oriented comprehension under long-context inputs. DuReader He et al. (2017) is a Chinese machine reading comprehension dataset with over human-annotated question-answer pairs from Baidu search logs. It covers open-domain QA with a range of answer formats. We use the English-translated version and evaluate factual correctness. The dataset is hosted at https://github.com/baidu/DuReader. Multi-News Fabbri et al. (2019) is a multi-document summarization dataset containing news articles from multiple sources clustered around the same event, with human-written summaries. It is accessible at https://github.com/Alex-Fabbri/Multi-News. This dataset evaluates the model’s ability to synthesize content across multiple documents and is especially useful for benchmarking performance on broad-context summarization.
@@ D  Limitations and Broader Impacts
Limitations. While MorphServe is practical and effective for dynamic LLM serving, several limitations remain. To support runtime layer swapping, MorphServe stores both full-precision and quantized variants in host memory. Although this increases memory usage, the overhead is typically under $2\times$ the model size and is well accommodated by modern LLM serving clusters. Future work may further reduce this cost by streaming layers from SSD to host memory on demand or directly fetching them from SSD via GPUDirect Storage (GDS) NVIDIA . MorphServe currently applies morphing at the transformer layer level. While effective, finer-grained adaptation, such as independently adjusting attention and MLP submodules, could unlock additional efficiency and precision flexibility. MorphServe reacts to system pressure in real time but does not anticipate upcoming surges. Integrating lightweight workload forecasting could enable proactive morphing decisions and further improve responsiveness under bursty traffic.